OpenStudy (anonymous):

The density of iron is 7.87 g/mL. What is the mass of a 3.00cm × 4.00 cm × 1.00 cm block of iron?

OpenStudy (jhannybean):

If \(\large D = \frac{M}{V}\) then \(\large \text{mass} = D\cdot V\) first calculate the volume. \(\large V=(3.00 \ cm)(4.00 \ cm)(1.00 \ cm) = ? \ cm^3 \) then multiply it to the density. \(\large D= 7.87 \frac{g}{mL}\)

OpenStudy (jhannybean):

And since \(\ 1 \ mL =1 \ cm^3\) The units cancel out giving you grams.

OpenStudy (jhannybean):

\[\large m= 12.00 \ mL \times 7.87 \frac{g}{mL} = (12.00)(7.87) \ g\]